At the movies concession stand, the cost for one popcorn and one drink id $5.75. The cost for 3 popcorns and 2 drinks is 14.75. what is the cost for popcorn?
Answer by stanbon(75887)

At the movies concession stand, the cost for one popcorn and one drink id $5.75. The cost for 3 popcorns and 2 drinks is 14.75. what is the cost for popcorn?
---------------------------------------
Equations:
p + d = 575 cents
3p+2d = 1475 cents
---------------------
Multiply thru the 1st eq. by 2 to get:
2p + 2d = 2*575
---
Subtract that from the 2nd eq. and solve for "p":
p = 1475-2*575 = 325 cents
---
popcorn = $3.75
==========================
Cheers,
Stan H.